### v2

@@ 3 Inconsistent Systems
Under consistent systems, and in the absence of round-off errors, the orthogonality and $\mathbf{A}$-conjugacy are the distinctive features of the CD methods. These properties play a key role in the development of the characteristically clean and concise algorithmic procedure of these methods. However, as long as the algorithm has not encountered an unlucky breakdown, these properties continue to hold even in inconsistent settings. The proof of Section 3 follows the standard line of reasoning in the consistent case, and hence is deferred to Appendix A; see Lemmas A.3 and A.1.
{property}
@@
Note that Section 3 does not require Algorithm 2 since these properties hold prior to a potential occurrence of an unlucky breakdown. Indeed, Algorithm 2 is only required when we need to ensure that Algorithms 1 and 2 can successfully generate all iterates up to termination.
@@ 3.1 Conjugate Gradient
Under an inconsistent system, there is a non-trivial, and a prior unknown, lower bound on the norm of the residual. Indeed, $\left\|\mathbf{r}\right\|^{2}=\left\|\mathbf{b}-\mathbf{Ax}\right\|^{2}=\|\mathbf{b}^{\perp}\|^{2}+\|\tilde{\mathbf{b}}-\mathbf{Ax}\|^{2}\geq\|\mathbf{b}^{\perp}\|^{2}$. Thus, termination criterion based on $\left\|\mathbf{r}\right\|$ is no longer practical. The alternative is to observe $\left\|\mathbf{Ar}\right\|\rightarrow 0$, as $\|\mathbf{Ar}\|=0$ indicates achieving a normal solution. However, we now show that CG is fundamentally unable to recover normal solutions in such settings. While it has been empirically observed that CG exhibits numerical instability under inconsistent systems, Theorem 3.1 provides a theoretical analysis of CG’s failure to converge in such settings.
Theorem 3.1 (CG’s Inability to Recover Normal Solutions).
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$ and $\mathbf{b}\notin\text{Null}(\mathbf{A})$. Under Algorithm 2, in Algorithm 1, we have $\mathbf{Ar}_{k}\neq\mathbf{0}$, for all $10\leq k\leq g-$ 1.
Proof 3.2.
We prove the claim by contradiction. Clearly, by the assumption on $\mathbf{b}$, $\mathbf{Ar}_{0}=\mathbf{Ab}\neq\mathbf{0}$. Suppose for some $11\leq k\leq g-$ 1, $\mathbf{A}\mathbf{r}_{k}=\mathbf{0}$. Using $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$, we get $\mathbf{A}\mathbf{p}_{k}=\beta_{k-1}\mathbf{A}\mathbf{p}_{k-1}$. By the assumption on $\mathbf{b}$, we have $\beta_{k-1}\neq 0$. Subsequently, from $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$, we get $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k}/\beta_{k-1}$, and by the inner-product with $\mathbf{p}_{k-1}$ and using Section 3-(b), it gives $\left\langle\mathbf{p}_{k-1},\mathbf{r}_{k-1}\right\rangle=0$. However, by Section 3-(b), we also have $\left\langle\mathbf{p}_{k-1},\mathbf{r}_{k-1}\right\rangle=\left\|\mathbf{r}_{k-1}\right\|^{2}\neq 0$, for $10\leq k\leq g-$ 1, which leads to a contradiction.
While Theorem 3.1 for inconsistent systems stands in sharp contrast to Fact 1 for consistent systems, it turns out that the last iterate of CG for inconsistent settings can still be characterized, in an analogous way to Fact 2 for consistent systems.
Lemma 3.3.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Algorithm 2, in Algorithm 1 we have $\mathbf{Ap}_{k-1}=\mathbf{0}$ if and only if $k=g$. Furthermore, $\mathbf{p}_{g-1}\neq\mathbf{0}$.
Proof 3.4.
Let $k=g$. By Section 3-(a) and Definition 1.2, we note that $\mathbf{Ap}_{g-1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ and
@@
In addition, from Section 3-(b), we have $\left\langle\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=0$, for all $11\leq i<g-$ 1. For $i=1$, $0=\left\langle\mathbf{p}_{0},\mathbf{Ap}_{g-1}\right\rangle=c_{0}\left\langle\mathbf{p}_{0},\mathbf{r}_{0}\right\rangle=c_{0}\|\mathbf{r}_{0}\|^{2}$, which since $\|\mathbf{r}_{0}\|\neq 0$ implies $c_{0}=0$. Similarly, for $12\leq i<g-$ 1,
@@
where the last equality follows from Section 3-(b). Since $\|\mathbf{r}_{i-1}\|\neq 0$, it follows that $c_{i-1}=0$. As a result, we must have $\mathbf{Ap}_{g-1}=c_{g-1}\mathbf{r}_{g-1}$. Now, if $c_{g-1}=0$, then we are done. Otherwise, assuming $\mathbf{Ap}_{g-1}\not=\mathbf{0}$, since $\mathbf{r}_{g-1}\neq\mathbf{0}$, we can, in principle, proceed to the next iteration by calculating $\alpha_{g-1}=\|\mathbf{r}_{g-1}\|^{2}/\left\langle\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle$. We first note that by Section 3-(b), we have
@@
Conversely, let $\mathbf{Ap}_{k-1}=\mathbf{0}$ for some $1\leq k\leq g$. Firstly, we show that $\mathbf{p}_{k-1}\not=\mathbf{0}$. Clearly, if $k=1$, we have $\mathbf{p}_{0}=\mathbf{b}\neq\mathbf{0}$. If $k\geq 2$, from $\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$ and Section 3-(b), it follows that
@@
As a result, since $\mathbf{p}_{k-1}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{p}_{k-1}\neq\mathbf{0}$, we can write $\mathbf{p}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i}\mathbf{b}$ and at least one $c_{j}$ must be non-zero. Hence, $\mathbf{0}=\mathbf{Ap}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i+1}\mathbf{b}$, which implies $\dim(\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}))=\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))$, and so we must have $k=g$.
Remark 3.5.
There is a subtlety in how the grade of $\mathbf{b}$ with respect to $\mathbf{A}$ is reflected in the total number of iterations before termination. If $\mathbf{b}\in\text{Range}(\mathbf{A})$, CG terminates when $k=g$ (Facts 1 and 2), while for $\mathbf{b}\notin\text{Range}(\mathbf{A})$, the termination occurs at $k=g-1$ (Lemma 3.3). However, in both cases, the total number of iterations is determined by the number of non-zero $\mathbf{b}$-relevant eigenvalues [^4] These are distinct eigenvalues whose corresponding eigenspace is not entirely orthogonal to $\mathbf{b}$; see the work 18 for more details of $\mathbf{A}$. Indeed, suppose $\mathbf{b}\notin\text{Range}(\mathbf{A})$ and consider its decomposition as in 2. Let $g$ and $\tilde{g}$ denote the grade of $\mathbf{b}$ and $\tilde{\mathbf{b}}$, respectively, with respect to $\mathbf{A}$. CG$(\mathbf{A},\tilde{\mathbf{b}})$ and CG$(\mathbf{A},\mathbf{b})$ terminate after $\tilde{g}$ and $g-1$ iterations, respectively. On the other hand, noting that $\tilde{\mathbf{b}}$ is orthogonal to the eigenspace corresponding to the zero eigenvalues of $\mathbf{A}$, for any eigenvector $\mathbf{v}$ of $\mathbf{A}$, we have
@@
That is, the number of $\mathbf{b}$-relevant eigenvalues is exactly one more than the number of $\tilde{\mathbf{b}}$-relevant ones. As a result, from Lemma 1 in the work of Liu and Roosta18, it follows that $\tilde{g}=g-1$, which implies that both situations result in precisely the same number of iterations.
Theorem 3.1 implies that, for inconsistent systems, CG’s output might not necessarily correspond to any solutions. Once $k=g-1$, by Lemma 3.3, $\alpha_{g-1}$ is no longer well-defined and the update direction $\mathbf{p}_{g-1}$ is not used. One may naïvely attempt to construct $\mathbf{x}_{g}=\mathbf{x}_{g-1}+\alpha\mathbf{p}_{g-1}$ for some suitable $\alpha$ in an effort to recover a normal solution. However, for any such update $\mathbf{x}_{g}=\mathbf{x}_{g-1}+\alpha\mathbf{p}_{g-1}$, we have $\mathbf{r}_{g}=\mathbf{r}_{g-1}$, and hence by Theorem 3.1, $\mathbf{x}_{g}$ cannot be a normal solution. Nonetheless, we now show that a modification to CG theoretically allows for the recovery of the pseudo-inverse solution in inconsistent settings.
To do this, we make two key observations. First, by Lemma B.3 in Appendix B, the Krylov subspace $\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ contains a normal solution. Denoting this normal solution as $\mathbf{x}^{\star}$, under Algorithm 2 and using Section 3-(a), we can write
@@
Lemma 3.6.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Algorithm 2, in Algorithm 1, we have
@@
Proof 3.7.
By Lemma B.1 in Appendix B, we know that $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{g-1}=a_{g-1}\mathbf{b}^{\perp}$, which by Lemma 3.3 gives $\mathbf{p}_{g-1}=a_{g-1}\mathbf{b}^{\perp}$. Furthermore, Lemma 3.3 gives $\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle=\left\langle\mathbf{p}_{g-1},\tilde{\mathbf{b}}\right\rangle+\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle=\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle$. Hence,
@@
Finally, using Section 3-(b) and the definition of $\beta_{k-1}$, we get $\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle=\|\mathbf{r}_{g-1}\|^{2}$.
@@
where the first equality uses Lemma 3.6, and the fourth equality employs Lemma 3.3 and the fact that in Algorithm 1, $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$. Now, for each $11\leq k\leq g-$ 1, we have
@@
which using Section 3-(b) gives
@@
$\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-1}\right\rangle$ $=\left\langle\mathbf{p}_{k-1},\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2}\right\rangle=\beta_{g-2}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-2}\right\rangle=\beta_{g-2}\beta_{g-3}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-3}\right\rangle=\ldots=\prod_{j=2}^{g-(k-1)}\beta_{g-j}\|\mathbf{p}_{k-1}\|^{2}=\frac{\|\mathbf{r}_{g-1}\|^{2}}{\|\mathbf{r}_{k-1}\|^{2}}\|\mathbf{p}_{k-1}\|^{2},$
@@
for $11\leq k\leq g-$ 1. Using these scalars, coupled with the particular choice $\gamma_{g-1}=0$, $\mathbf{x}^{\star}$ in 7 now amounts to a normal solution. To find the pseudo-inverse solution, we note that, Lemma B.1 implies
@@
where $c=\sum_{k=1}^{g}\eta_{k-1}a_{k-1}$. However, using Lemma 3.3,
@@
Note that the left-hand side can also be regarded as the orthogonal projection of $\mathbf{x}^{\star}$ onto the orthogonal complement of $\text{Range}(\mathbf{p}_{g-1})$. This leads us to Algorithms 3 and 3.8.
@@
Theorem 3.8 (Pseudo-inverse Solution with Algorithm 3).
Under Algorithm 2, in Algorithm 3 we have $\mathbf{x}^{+}=\mathbf{A}^{\dagger}\mathbf{b}$.
Remark 3.9.
We highlight that while, theoretically, Algorithm 3 provides a means to extract the pseudo-inverse solution at termination, the extreme numerical instability of CG for inconsistent systems causes the theory to fall short of practical observations. The instability of CG under inconsistent systems is widely observed in many applications and also documented in the works 24, 26; see also the experiments in Section 4.

### v4

@@ 3 Inconsistent Systems
Under consistent systems, and in the absence of round-off errors, the orthogonality and $\mathbf{A}$-conjugacy are the distinctive features of the CD methods. These properties play a key role in the development of the characteristically clean and concise algorithmic procedure of these methods. However, as long as the algorithm has not encountered an unlucky breakdown, these properties continue to hold even in inconsistent settings. The proof of Property 1 follows the standard line of reasoning in the consistent case, and hence is deferred to Appendix A; see Lemmas 10 and 9.
Property 1.
@@
Note that Property 1 does not require Assumption 1 since these properties hold prior to a potential occurrence of an unlucky breakdown. Indeed, Assumption 1 is only required when we need to ensure that Algorithms 1 and 2 can successfully generate all iterates up to termination.
@@ 3.1 Conjugate Gradient
Under an inconsistent system, there is a non-trivial, and a prior unknown, lower bound on the norm of the residual. Indeed, $\left\|\mathbf{r}\right\|^{2}=\left\|\mathbf{b}-\mathbf{Ax}\right\|^{2}=\|\mathbf{b}^{\perp}\|^{2}+\|\tilde{\mathbf{b}}-\mathbf{Ax}\|^{2}\geq\|\mathbf{b}^{\perp}\|^{2}$. Thus, termination criterion based on $\left\|\mathbf{r}\right\|$ is no longer practical. The alternative is to observe $\left\|\mathbf{Ar}\right\|\rightarrow 0$, as $\|\mathbf{Ar}\|=0$ indicates achieving a normal solution. However, we now show that CG is fundamentally unable to recover normal solutions in such settings. While it has been empirically observed that CG exhibits numerical instability under inconsistent systems, Theorem 1 provides a theoretical analysis of CG’s failure to converge in such settings.
Theorem 1 (CG’s Inability to Recover Normal Solutions).
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$ and $\mathbf{b}\notin\text{Null}(\mathbf{A})$. Under Assumption 1, in Algorithm 1, we have $\mathbf{Ar}_{k}\neq\mathbf{0}$, for all $10\leq k\leq g-$ 1.
Proof.
We prove the claim by contradiction. Clearly, by the assumption on $\mathbf{b}$, $\mathbf{Ar}_{0}=\mathbf{Ab}\neq\mathbf{0}$. Suppose for some $11\leq k\leq g-$ 1, $\mathbf{A}\mathbf{r}_{k}=\mathbf{0}$. Using $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$, we get $\mathbf{A}\mathbf{p}_{k}=\beta_{k-1}\mathbf{A}\mathbf{p}_{k-1}$. By the assumption on $\mathbf{b}$, we have $\beta_{k-1}\neq 0$. Subsequently, from $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$, we get $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k}/\beta_{k-1}$, and by the inner-product with $\mathbf{p}_{k-1}$ and using Property 1-(b), it gives $\left\langle\mathbf{p}_{k-1},\mathbf{r}_{k-1}\right\rangle=0$. However, by Property 1-(b), we also have $\left\langle\mathbf{p}_{k-1},\mathbf{r}_{k-1}\right\rangle=\left\|\mathbf{r}_{k-1}\right\|^{2}\neq 0$, for $10\leq k\leq g-$ 1, which leads to a contradiction. ∎
Remark 4.
CG’s inability to recover a solution of the normal equation can be better understood through the framework from which it is derived. Specifically, CG is a particular algorithmic implementation of the Orthogonal-Residual (OR) framework [15, 4], which applies to square, but not necessarily symmetric, matrices. When restricted to PD matrices, the OR framework can be algorithmically implemented as Algorithm 1. However, in Appendix C, we demonstrate that any algorithm derived from the OR framework cannot converge to a normal solution for symmetric but inconsistent systems (see Theorem 6). Since CG is derived from OR, it inherits this “non-convergent” property.
While Theorem 1 for inconsistent systems stands in sharp contrast to Fact 1 for consistent systems, it turns out that the last iterate of CG for inconsistent settings can still be characterized, in an analogous way to Fact 2 for consistent systems.
Lemma 1.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 1 we have $\mathbf{Ap}_{k-1}=\mathbf{0}$ if and only if $k=g$. Furthermore, $\mathbf{p}_{g-1}\neq\mathbf{0}$.
Proof.
Let $k=g$. By Property 1-(a) and Definition 2, we note that $\mathbf{Ap}_{g-1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ and
@@
In addition, from Property 1-(b), we have $\left\langle\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=0$, for all $11\leq i<g-$ 1. For $i=1$, $0=\left\langle\mathbf{p}_{0},\mathbf{Ap}_{g-1}\right\rangle=c_{0}\left\langle\mathbf{p}_{0},\mathbf{r}_{0}\right\rangle=c_{0}\|\mathbf{r}_{0}\|^{2}$, which since $\|\mathbf{r}_{0}\|\neq 0$ implies $c_{0}=0$. Similarly, for $12\leq i<g-$ 1,
@@
where the last equality follows from Property 1-(b). Since $\|\mathbf{r}_{i-1}\|\neq 0$, it follows that $c_{i-1}=0$. As a result, we must have $\mathbf{Ap}_{g-1}=c_{g-1}\mathbf{r}_{g-1}$. Now, if $c_{g-1}=0$, then we are done. Otherwise, assuming $\mathbf{Ap}_{g-1}\not=\mathbf{0}$, since $\mathbf{r}_{g-1}\neq\mathbf{0}$, we can, in principle, proceed to the next iteration by calculating $\alpha_{g-1}=\|\mathbf{r}_{g-1}\|^{2}/\left\langle\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle$. We first note that by Property 1-(b), we have
@@
Conversely, let $\mathbf{Ap}_{k-1}=\mathbf{0}$ for some $1\leq k\leq g$. Firstly, we show that $\mathbf{p}_{k-1}\not=\mathbf{0}$. Clearly, if $k=1$, we have $\mathbf{p}_{0}=\mathbf{b}\neq\mathbf{0}$. If $k\geq 2$, from $\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$ and Property 1-(b), it follows that
@@
As a result, since $\mathbf{p}_{k-1}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{p}_{k-1}\neq\mathbf{0}$, we can write $\mathbf{p}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i}\mathbf{b}$ and at least one $c_{j}$ must be non-zero. Hence, $\mathbf{0}=\mathbf{Ap}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i+1}\mathbf{b}$, which implies $\dim(\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}))=\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))$, and so we must have $k=g$. ∎
Remark 5.
There is a subtlety in how the grade of $\mathbf{b}$ with respect to $\mathbf{A}$ is reflected in the total number of iterations before termination. If $\mathbf{b}\in\text{Range}(\mathbf{A})$, CG terminates when $k=g$ (Facts 1 and 2), while for $\mathbf{b}\notin\text{Range}(\mathbf{A})$, the termination occurs at $k=g-1$ (Lemma 1). However, in both cases, the total number of iterations is determined by the number of non-zero $\mathbf{b}$-relevant eigenvalues [^7] These are distinct eigenvalues whose corresponding eigenspace is not entirely orthogonal to $\mathbf{b}$; see the work [27] for more details of $\mathbf{A}$. Indeed, suppose $\mathbf{b}\notin\text{Range}(\mathbf{A})$ and consider its decomposition as in 2. Let $g$ and $\tilde{g}$ denote the grade of $\mathbf{b}$ and $\tilde{\mathbf{b}}$, respectively, with respect to $\mathbf{A}$. CG$(\mathbf{A},\tilde{\mathbf{b}})$ and CG$(\mathbf{A},\mathbf{b})$ terminate after $\tilde{g}$ and $g-1$ iterations, respectively. On the other hand, noting that $\tilde{\mathbf{b}}$ is orthogonal to the eigenspace corresponding to the zero eigenvalues of $\mathbf{A}$, for any eigenvector $\mathbf{v}$ of $\mathbf{A}$, we have
@@
That is, the number of $\mathbf{b}$-relevant eigenvalues is exactly one more than the number of $\tilde{\mathbf{b}}$-relevant ones. As a result, from Lemma 1 in the work of Liu and Roosta[27], it follows that $\tilde{g}=g-1$, which implies that both situations result in precisely the same number of iterations.
Theorem 1 implies that, for inconsistent systems, CG’s output might not necessarily correspond to any solutions. Once $k=g-1$, by Lemma 1, $\alpha_{g-1}$ is no longer well-defined and the update direction $\mathbf{p}_{g-1}$ is not used. One may naïvely attempt to construct $\mathbf{x}_{g}=\mathbf{x}_{g-1}+\alpha\mathbf{p}_{g-1}$ for some suitable $\alpha$ in an effort to recover a normal solution. However, for any such update $\mathbf{x}_{g}=\mathbf{x}_{g-1}+\alpha\mathbf{p}_{g-1}$, we have $\mathbf{r}_{g}=\mathbf{r}_{g-1}$, and hence by Theorem 1, $\mathbf{x}_{g}$ cannot be a normal solution. Nonetheless, we now show that a modification to CG theoretically allows for the recovery of the pseudo-inverse solution in inconsistent settings.
To do this, we make two key observations. First, by Lemma 15 in Appendix B, the Krylov subspace $\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ contains a normal solution. Denoting this normal solution as $\mathbf{x}^{\star}$, under Assumption 1 and using Property 1-(a), we can write
@@
Lemma 2.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 1, we have
@@
Proof.
By Lemma 12 in Appendix B, we know that $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{g-1}=a_{g-1}\mathbf{b}^{\perp}$, which by Lemma 1 gives $\mathbf{p}_{g-1}=a_{g-1}\mathbf{b}^{\perp}$. Furthermore, Lemma 1 gives $\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle=\left\langle\mathbf{p}_{g-1},\tilde{\mathbf{b}}\right\rangle+\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle=\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle$. Hence,
@@
Finally, using Property 1-(b) and the definition of $\beta_{k-1}$, we get $\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle=\|\mathbf{r}_{g-1}\|^{2}$. ∎
@@
where the first equality uses Lemma 2, and the fourth equality employs Lemma 1 and the fact that in Algorithm 1, $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$. Now, for each $11\leq k\leq g-$ 1, we have
@@
which using Property 1-(b) gives
@@
$\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-1}\right\rangle$ $=\left\langle\mathbf{p}_{k-1},\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2}\right\rangle=\beta_{g-2}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-2}\right\rangle=\beta_{g-2}\beta_{g-3}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-3}\right\rangle=\ldots$
$=\prod_{j=2}^{g-(k-1)}\beta_{g-j}\|\mathbf{p}_{k-1}\|^{2}=\frac{\|\mathbf{r}_{g-1}\|^{2}}{\|\mathbf{r}_{k-1}\|^{2}}\|\mathbf{p}_{k-1}\|^{2},$
@@
for $11\leq k\leq g-$ 1. Using these scalars, coupled with the particular choice $\gamma_{g-1}=0$, $\mathbf{x}^{\star}$ in 7 now amounts to a normal solution. To find the pseudo-inverse solution, we note that, Lemma 12 implies
@@
where $c=\sum_{k=1}^{g}\eta_{k-1}a_{k-1}$. However, using Lemma 1,
@@
Note that the left-hand side can also be regarded as the orthogonal projection of $\mathbf{x}^{\star}$ onto the orthogonal complement of $\text{Range}(\mathbf{p}_{g-1})$. This leads us to Algorithms 3 and 2.
@@
Theorem 2 (Pseudo-inverse Solution with Algorithm 3).
Under Assumption 1, in Algorithm 3 we have $\mathbf{x}^{+}=\mathbf{A}^{\dagger}\mathbf{b}$.
Remark 6.
We highlight that while, theoretically, Algorithm 3 provides a means to extract the pseudo-inverse solution at termination, the extreme numerical instability of CG for inconsistent systems causes the theory to fall short of practical observations. The instability of CG under inconsistent systems is widely observed in many applications and also documented in the works [23, 35]; see also the experiments in Section 4.